# Incorporating Flexible Image Conditioning into Text-to-Video Diffusion Models without Training

### v1

@@ 5.3 Visualization for Qualitative Comparison
Table 2: Ablation study of components in our method. $\downarrow$ means a lower score in this metric indicates a better performance. The best results are highlighted with boldface. The orange row refers to our complete method. Each component is removed from the complete method separately.
@@ 5.4 Ablation Study
In our method, there are two components (frame replacement and random patch swapping) to incorporate visual conditioning into T2V diffusion models. We evaluate the contribution of each component by ablating them from our method separately. The results are presented in Tab. 2. After removing frame replacement, the performance lags behind the full model by a small margin in all tasks. The result indicates that random patch swapping (RPS) can work alone to effectively merge visual guidance into T2V diffusion process. It also validates our hypothesis that sufficient visual features can be transferred through a portion of local patches at each denoising step for video generation. On the contrary, if we remove RPS and only implement frame replacement strategy for extra image conditioning, a significant performance drop is observed across all metrics and tasks. Though frame replacement introduces image features into the temporal layers of UNet, the model estimates the holistic noise added to all video frames (16 frames in our experimental setting) as an entirety. Hence, the condition image features are likely to be ignored because of the small number of images against the large number of video frames. We observe that the videos generated without RPS greatly deviate from the condition images, thus resulting in a remarkable decrease in the performance. More analysis of our methods is provided in the supplementary.
Figure 5: Generatlization to videos without humans. The image with blue edges is the condition image, and the images with green edges are generated by our method.
5.5 Generalization to Videos without Humans
We further evaluate the generalization capability of our method to videos without humans. As demonstrated in Fig. 5, we prompt our method to synthesize a video with a panda as the subject. The output video well aligns with the textual instruction and preserves the visual features in the condition image, validating the robustness of our method. More examples are illustrated in the supplementary.
@@ 6 Conclusion
Appendices and Supplementary Material
This is the supplementary material for the paper titled “Incorporating Flexible Image Conditioning into Text-to-Video Diffusion Models without Training”. We organize the content as follows:
@@
B.1 – Comparison of Inference Speed between FlexTI2V and Baselines
B.2 – Analysis of Dynamic Control Strategy
B.3 – Visualization for Ablation Study
B.4 – More Visualization of Generated Videos without Humans
B.5 – Additional Demonstration of Our Method
B.6 – Implementation of Our Method on Wan2.1-T2V
B.7 – Failure Cases
@@
E – Potential Societal Impacts of Our Research
F – Code and Data Release
Appendix A Algorithm Formulation
We formulate our FlexTI2V approach in Alg. 1. The function $Repeat+(\tilde{x}_{T}^{(0)},M)$ at step 3 duplicates the inverted condition image representation for $M$ times as the initialization for video frame denoising. If there are more than one condition images, we find using different image makes trivial difference in the generation performance. We use the first condition image in our method. The function $BoundIndex+(m,\bm{p})$ at step 8 returns the indices of the two closest condition images before and after $m$-th video frame respectively (i.e., bounded index). If all condition images are before or after this video frame, the function returns only one index of the closest image.
Appendix B Additional Experiment Resutls
B.1 Comparison of Inference Speed between FlexTI2V and Baselines
We evaluate the average inference time of our method and the baseline models for generating a single 16-frame video in Tab. 3. The test is conducted on one NVIDIA H100 GPU. TI2V-Zero synthesizes a single video frame in one denoising process, and relies on multiple resample steps to improve the performance. The inference time of TI2V-Zero thus is significantly longer than other approaches. TRF conducts denoising processes conditioned on two bounded images separately and then merges the features at each step, resulting in longer inference time than DynamiCrafter. In contrast, our method requires only 2.71 seconds for 20 DDIM denosing steps for video generation, which is much lower than DynamiCrafter and TRF. The great gap validates the high inference efficiency of our method.
Algorithm 1 FlexTI2V
@@ B.1 Comparison of Inference Speed between FlexTI2V and Baselines
3:$\tilde{\bm{x}}_{0}\leftarrow\mathcal{E}(\bm{x})$
4:$\tilde{\bm{x}}_{T}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{T}}\;\tilde{\bm{x}}_{0},\,\sqrt{1-\bar{\alpha}_{t}}\;\bm{I}\right)$
5:$\bm{z}_{T}\leftarrow\verb+Repeat+(\tilde{x}_{T}^{(0)},\ M)$triangleright Noise Initialization
@@
7:    $\tilde{\bm{x}}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\;\tilde{\bm{x}}_{0},\sqrt{1-\bar{\alpha}_{t}}\;\bm{I}\right)$triangleright Image Inversion
8:    $\hat{\bm{z}}_{t}\leftarrow\verb+FrameReplace+(\tilde{\bm{x}}_{t},\ \bm{z}_{t},\ \bm{p})$triangleright Frame Replacement, Eq. 3 in Main Paper
@@
10:         $\bm{n}_{bound}\leftarrow\verb+BoundIndex+(m,\bm{p})$triangleright Indices of Bounded Condition Image
@@
12:             $\tilde{t}\leftarrow t_{0}-\delta_{2}\cdot|m-p^{(n)}|$triangleright Dynamic Control in Swapping Steps
@@
14:                 $P(m,n,t)\leftarrow P_{0}-\delta_{1}\cdot|m-p^{(n)}|$triangleright Dynamic Control in Swapping Percentage
@@
18:             Generate random binary mask $\mathcal{M}_{t}^{(m,n)}$ with element “$1$” accounting for $P(m,n,t)$.
@@
25:    $\bm{z}_{t-1}\leftarrow\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right)+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)$triangleright Deterministic Sampling
@@
Figure 6: Demonstration of ablation study. The images with blue edges are condition images and images with green edges are generated video frames. More explanations are provided in Sec. B.2 and Sec. B.3.
@@
Figure 7: Additional examples of our method on videos without humans. The images with blue edges are condition images and images with green edges are generated video frames. Please see Sec. B.4 for more details.
B.2 Analysis of Dynamic Control Strategy
@@
We remove the dynamic control strategy from our method and use a fixed percentage and steps for random patch swapping. The results are illustrated in the last row of Fig. 6. The motion in the generated video frames is trivial without dynamic control. The reason is because the constraints of visual conditioning are overly strong for all frames. The result validates the necessity of the dynamic control strategy.
B.3 Visualization for Ablation Study
@@
Besides ablation of dynamic control (see Sec. B.2), we also show additional visualization for quantitative ablation study of patch replacement and random patch swapping in Fig. 6. After removing patch replacement, the output video is still decent with minor deviation from the condition image. However, without random patch swapping, the generated video fails to follow the visual condition, thus leading to a poor performance in main paper Tab. 2.
DynamiCrafter [54] TI2V-Zero [37] TRF [7] FlexTI2V (Ours)
Inference Time (s) 9.45 275.14 60.05 2.71
Table 3: Comparison of model efficiency in inference. The best results are highlighted with boldface.
B.4 More Visualization of Generated Videos without Humans
@@
The videos in UCF-101 are mainly about human actions. To assess the generalization performance of our approach, we illustrate the demonstration of FlexTI2V synthesizing videos without humans in main paper Sec. 5.5 and Fig. 5. We supplement extra examples in Fig. 7. Our approach seamlessly generalizes to the motion of objects showing strong generalization capability.
@@
Figure 8: Additional demonstration of videos synthesized by FlexTI2V in classic TI2V settings. The images with blue and pink edges are condition imags, and images with green edges are generated video frames. $N$ and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
@@
Figure 9: Additional demonstration of videos synthesized by FlexTI2V in new TI2V settings. The images with blue, pink and orange edges are condition imags, and images with green edges are generated video frames. $N$ and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
B.5 Additional Demonstration of Our Method
@@
We illustrate more videos synthesized by our method in classic TI2V generation settings (see Fig. 8) and new settings (see Fig. 9). Our approach is able to condition the pretrained T2V diffusion foundation model on any number of images with flexible positions.
Figure 10: Generated videos of our method implemented to Wan2.1-T2V. The images with blue edges are condition images and images with green edges are generated video frames. Please refer to Sec. B.6 for more explanation.
B.6 Implementation of Our Method on Wan2.1-T2V
We use ModelScopeT2V [52] as the T2V diffusion foundation model in our experiments for a fair comparison with previous approaches. We also implement our method to another state-of-the-art T2V foundation model – Wan2.1-T2V [50]. The results in Fig. 10 suggest that our method can be applied to to other diffusion foundation models, showing strong generalization capability.
@@
Figure 11: Typical failure cases of our approach. The images with blue and pink edges are condition images and images with green edges are generated video frames. Please refer to Sec. B.7 for more analysis.
B.7 Failure Cases
@@
We demonstrate some failure cases of our method in Fig. 11 to show the boundary of the capability. We find it is challenging for our method to synthesize smooth camera viewpoint change in output videos (first example). If there is a big difference in viewpoints of two condition images, our method is very likely to treat the two images as irrelevant images and generates a hard transition without a reasonable motion. In addition, we also observe watermarks in a few output videos (second example). This is an inherent bias of the foundation model, which is inherited by our method. Hence, the performance of our method is still limited by the pretrained T2V foundation model.
Appendix C Implementation Details

### v2

@@ 5.2 Quantitative Comparison
Table 3: Ablation study of components in our method. $\downarrow$ means a lower score in this metric indicates a better performance. The orange row refers to our complete method. Each component is removed from the complete method separately.
Multi-Image Conditioning In terms of video frame interpolation, DynamiCrafter still lags behind TRF and FlexTI2V due to domain shift. In contrast, TRF shows better performance for this task (measured by FVD) relying on the strong generalization capability of stable video diffusion [blattmann2023stable], which further validates the advantage of training-tree method in controllable video generation. Though TRF achieves a high score in CLIP-F and CLIP-I, it still has a limited performance in video-text alignment (tFVD and CLIP-T) due to lack of semantic control by textual prompts. In contrast, our method outperforms both competitors in all the metrics. The significant improvement further validates the flexibility of our method to multi-image conditioning.
User Study Besides automatic metrics. We also conduct user study for a thorough evaluation. We sample 100 instances from our test set for each task, with each instance evaluated by five raters. Each rater is asked to select the video with best alignment with the textual prompt and condition images. The human preference percentage of each method is presented in Tab. 1. FlexTI2V is selected as the best in the highest percentage of samples compared with other competitors. The results suggest that the videos synthesized by our method aligns better with human’s subjective perception.
Inference Efficiency We evaluate the average inference time of our method and the baseline models for generating a single 16-frame video in Tab. 2. The test is conducted on one NVIDIA H100 GPU. TI2V-Zero synthesizes video frames one by one, and relies on multiple resample steps to improve the performance. The inference time thus is significantly longer than other approaches. TRF conducts denoising processes conditioned on two bounded images separately and then merges the features at each step, resulting in longer inference time than DynamiCrafter. In contrast, our method requires only 2.71 seconds for 20 DDIM denosing steps, which is much lower than DynamiCrafter and TRF. The great gap validates the high efficiency of our method.
5.3 Visualization for Qualitative Comparison
Besides quantitative evaluation, we also visualize the videos synthesized by our method and other baselines for a qualitative comparison. As illustrated in Fig. 4, we observe some distortion in the videos generated by DynamiCrafter (e.g., the person in image animation and the knife in video frame interpolation), which may be the explanation for the low performance in Tab. 1. TRF is good at preserving visual features of the two bounded images. However, the video frames synthesized by TRF are blurry around the knife. We speculate that TRF is initially proposed for smooth motion generation and camera viewpoint transition. It may be weak in modeling object state change, such as the onion being cut into small pieces in this example. TI2V-Zero can generate videos with large motion because of the autoregressive generation mechanism. However, the cumulative mistakes weaken the controllability of visual conditions, leading to a low consistency in the frames far from the condition image. In contrast, our method can follow the textual prompts and visual conditions faithfully, and synthesize smooth motions in all TI2V settings. Please refer to Secs. B.6 and B.7 in the supplementary for more visualization of our method.
5.4 Generalization to Different Foundation Models
We use ModelScopeT2V [wang2023modelscope] as the foundation model in previous experiments simply for a fair comparison with prior methods. We also implement our method to another SOTA T2V foundation model – Wan2.1-T2V [wang2025wan] using transformer-based architectures. As shown in Fig. 5, our method can also incorporate additional image conditioning to Wan2.1 and synthesize videos with high fidelity and smooth motions, showing strong generalization capability to different T2V models. The quantitative comparison with Wan2.1 model family is demonstrated in Supp. Sec. B.1.
@@ 5.5 Ablation Study
In our method, there are two components (frame replacement and random patch swapping) to incorporate visual conditioning into T2V diffusion models. We evaluate the contribution of each component by ablating them from our method separately. The results are presented in Tab. 3. After removing frame replacement, the performance lags behind the full model by a small margin in all tasks. The result indicates that random patch swapping (RPS) can work alone to effectively merge visual guidance into T2V diffusion process. It also validates our hypothesis that sufficient visual features can be transferred through a portion of local patches at each denoising step for video generation. On the contrary, if we remove RPS and only implement frame replacement strategy for extra image conditioning, a significant performance drop is observed across all metrics and tasks. Though frame replacement introduces image features into the temporal layers, the model estimates the holistic noise added to all video frames (16 frames in our experimental setting) as an entirety. Hence, the condition image features are likely to be ignored because of the small number of images against the large number of video frames. We observe that the videos generated without RPS greatly deviate from the condition images, thus resulting in a remarkable decrease in the performance. More ablation and hyperparameter analysis are provided in Supp. Secs. B.3, B.4 and B.5.
@@ 6 Conclusion
[…]
Supplementary Material
This is the supplementary material for the publication titled “Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning”. We organize the content as follows:
@@
B.1 – Quantitative Comparison with Wan2.1 Model Family
B.2 – Quantitative Results on Synthetic Dataset
B.3 – Analysis of Dynamic Control Strategy
B.4 – Visualization for Ablation Study
B.5 – Ablation of Hyperparameters
B.6 – Generated Videos without Humans
B.7 – Additional Demonstration of Our Method
B.8 – Failure Cases
@@
E – Code and Data Release
Algorithm A FlexTI2V
@@
[…]
@@
7:  $\tilde{\bm{x}}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\;\tilde{\bm{x}}_{0},\ (1-\bar{\alpha}_{t})\bm{I}\right)$ $\triangleright$ Image Inversion
8:  $\hat{\bm{z}}_{t}\leftarrow\verb+FrameReplace+(\tilde{\bm{x}}_{t},\ \bm{z}_{t},\ \bm{p})$ $\triangleright$ Frame Replacement, Eq.(3) in Main Paper
@@
10:    $\bm{n}_{bound}\leftarrow\verb+BoundIndex+(m,\bm{p})$ $\triangleright$ Indices of Bounded Condition Image
@@
12:     $\tilde{t}\leftarrow t_{0}-\delta_{2}\cdot|m-p^{(n)}|$ $\triangleright$ Dynamic Control in Swapping Steps
@@
14:      $P(m,n,t)\leftarrow P_{0}-\delta_{1}\cdot|m-p^{(n)}|$ $\triangleright$ Dynamic Control in Swapping Percentage
@@
18:     Generate random binary mask $\mathcal{M}_{t}^{(m,n)}$ with element “11” accounting for $P(m,n,t)$.
@@
25:  $\bm{z}_{t-1}\leftarrow\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right)+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)$ $\triangleright$ Deterministic Sampling
@@
A Algorithm Formulation
We formulate our FlexTI2V approach in Alg. A. The function $\verb+Repeat+(\tilde{x}_{T}^{(0)},M)$ at step 3 duplicates the inverted condition image representation for $M$ times as the initialization for video frame denoising. If there are more than one condition images, we find using different image makes trivial difference in the generation performance. We use the first condition image in our method. The function $\verb+BoundIndex+(m,\bm{p})$ at step 8 returns the indices of the two closest condition images before and after $m$-th video frame respectively (i.e., bounded index). If all condition images are before or after this video frame, the function returns only one index of the closest image.
B Additional Experiment Resutls
B.1 Quantitative Comparison with Wan2.1 Model Family
We additionally provide quantitative comparison with Wan2.1-I2V in image animation and Wan2.1-FLF2V in video frame interpolation on UCF-101 [soomro2012ucf101]. As shown in Tab. A and Tab. B Our method can be successfully applied to the SOTA T2V models and achieve a better performance compared with the finetuned counterpart. The possible reason is that it’s still challenging for training-based methods to generalize to out-of-distribution data (e.g., different aspect ratios and resolutions). This also shows the necessity of training-free methods like our work.
[…]
Wan2.1-I2V [wang2025wan] 141.22 31.01 82.33
Wan2.1-T2V + Our Method 101.74 30.98 85.87
Table A: Comparison with Wan2.1-I2V in image animation. The orange column refers to our method.
[…]
Wan2.1-FLF2V [github2025wan] 112.22 31.47 85.13
Wan2.1-T2V + Our Method 90.82 31.50 86.04
Table B: Comparison with Wan2.1-FLF2V in video interpolation. The orange column refers to our method.
Methods Animation Interpolation
CLIP-T CLIP-I CLIP-T CLIP-I
DynamiCrafter [xing2024dynamicrafter] 24.33 79.53 25.44 75.50
TRF [feng2024explorative] - - 25.92 81.38
TI2V-Zero [ni2024ti2v] 24.87 81.76 - -
FlexTI2V 26.83 82.31 27.29 83.54
Table C: Results on the synthetic dataset. The orange column refers to our method. Refer to Sec. B.2 for more descriptions.
B.2 Quantitative Results on Synthetic Dataset
In the main paper, we follow the previous work [ni2024ti2v] to conduct evaluation on UCF-101 for a fair comparison. To make the testing data more diverse, we utilize the data curation pipeline of [ni2024ti2v] to synthesize an OPEN dataset by Stable Video Diffusion using 10 prompts:
    “A mesmerizing display of the northern lights in the Arctic.”
    “A bustling street market in Marrakech with colorful textiles and spices.”
    “A futuristic cityscape with holographic advertisements and flying cars.”
    “A romantic gondola ride through the canals of Venice at sunset.”
    “A group of friends on a road trip, singing along to their favorite songs.”
    “A serene mountain cabin covered in a fresh blanket of snow.”
    “A thrilling skateboarder performing tricks in a skate park.”
    “A bustling night market in Bangkok with street food vendors and live music.”
    “A high-speed bullet train racing through a scenic countryside.”
    “A group of explorers uncovering the mysteries of an ancient temple in the jungle.”
Different from UCF-101, the 10 prompts are mainly about non-human motions and object state changes. The synthetic data also guarantees the base T2V model is not trained on it. We generate 22 videos for each prompt, totaling 220 testing samples.
As shown in Tab. C, our method still prominently outperforms all the baseline approaches in CLIP-T (text-video alignment) and CLIP-I (image-video similarity). The results validate that our method is also superior in modeling different types of motions and videos in the TI2V task.
@@ B.2 Quantitative Results on Synthetic Dataset
Figure A: Demonstration of ablation study. The images with blue edges are condition images and images with green edges are generated video frames. More explanations are provided in Sec. B.3 and Sec. B.4.
@@
Figure B: Additional examples of our method on videos without humans. The images with blue edges are condition images and images with green edges are generated video frames. Please see Sec. B.6 for more details.
B.3 Analysis of Dynamic Control Strategy
@@
We remove the dynamic control strategy from our method and use a fixed percentage and steps for random patch swapping. The results are illustrated in the last row of Fig. A. The motion in the generated video frames is trivial without dynamic control. The reason is because the constraints of visual conditioning are overly strong for all frames. The result validates the necessity of the dynamic control strategy.
B.4 Visualization for Ablation Study
@@
Besides ablation of dynamic control (see Sec. B.3), we also show additional visualization for quantitative ablation study of patch replacement and random patch swapping in Fig. A. After removing patch replacement, the output video is still decent with minor deviation from the condition image. However, without random patch swapping, the generated video fails to follow the visual condition, thus leading to a poor performance in main paper Tab. 3.
B.5 Analysis of Hyperparameters
In our experiments, the optimal hyperparameters are selected empirically. Below are the ablation studies for $\delta_{1}$, $\delta_{2}$, $P_{0}$ and $t_{0}$ on the image animation task. We change only one hyperparameter in each table.
As shown in Tab. D, if we use a large $\delta_{1}$ and $\delta_{2}$, or small $P_{0}$ and $t_{0}$, the image conditioning is weaker than the optimal hyperparameters so that the generated video can not accurately follow the images. Likewise, if $\delta_{1}$ and $\delta_{2}$ are small, or $P_{0}$ and $t_{0}$ are large, the image conditioning is too strong so that the output videos are nearly frozen. To achieve an optimal balance, we select the best hyperparameter combination based on the ablation study.
$\delta_{1}$ $4\times 10^{-3}$ $\bm{5\times 10^{-3}}$ $6\times 10^{-3}$ $7\times 10^{-3}$
[…]
$\delta_{2}$ $0.2$ $\bm{0.3}$ $0.4$ $0.5$
[…]
   $P_{0}$    $0.2$    $\bm{0.3}$    $0.4$
[…]
   $t_{0}$    77    $10$    $13$
[…]
Table D: Hyperparameter analysis. Please refer to Sec. B.5 for more details.
B.6 Generated Videos without Humans
@@
The videos in UCF-101 are mainly about human actions. To assess the generalization performance of our approach, we illustrate the demonstration of FlexTI2V synthesizing videos without humans in main paper Fig. 5. We supplement extra examples in Fig. B. Our approach seamlessly generalizes to the motion of objects showing strong generalization capability.
@@
Figure C: Additional demonstration of videos synthesized by FlexTI2V in classic TI2V settings. The images with blue and pink edges are condition imags, and images with green edges are generated video frames. NN and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
@@
Figure D: Additional demonstration of videos synthesized by FlexTI2V in new TI2V settings. The images with blue, pink and orange edges are condition imags, and images with green edges are generated video frames. NN and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
B.7 Additional Demonstration of Our Method
@@
We illustrate more videos synthesized by our method in classic TI2V generation settings (see Fig. C) and new settings (see Fig. D). Our approach is able to condition the pretrained T2V diffusion foundation model on any number of images with flexible positions.
@@
Figure E: Typical failure cases of our approach. The images with blue and pink edges are condition images and images with green edges are generated video frames. Please refer to Sec. B.8 for more analysis.
B.8 Failure Cases
@@
We demonstrate some failure cases of our method in Fig. E to show the boundary of the capability. We find it is challenging for our method to synthesize smooth camera viewpoint change in output videos (first example). If there is a big difference in viewpoints of two condition images, our method is very likely to treat the two images as irrelevant images and generates a hard transition without a reasonable motion. In addition, we also observe watermarks in a few output videos (second example). This is an inherent bias of the foundation model, which is inherited by our method. Hence, the performance of our method is still limited by the pretrained T2V foundation model.
C Implementation Details